OpenStudy (anonymous):

1)When mass and radius are held constant, what is the relationship between velocity and force? Explain and write as a proportionality.(v ∝F or v ∝1/F) 2)When mass and force are held constant, what is the relationship between velocity and radius? Explain and write as proportionality. 3)When force and radius are held constant, what is the relationship between velocity and mass? Answer in a sentence, and as proportionality. 4)Write a proportionality relating velocity to force, mass and radius. Rewrite it so we can see how force depends on other quantities.

OpenStudy (anonymous):

listen we knw F=mv^2/r now u are given v and r is constant so, \[\sqrt{F} \alpha Velocity\]

OpenStudy (anonymous):

now u are said mass and force are constant u ll get the ans by above formula again that \[V \alpha \sqrt{1/r}\]

OpenStudy (anonymous):

now force and radius are constant \[V \alpha \sqrt{1/m}\] i guess last one u can do yourself :)

OpenStudy (anonymous):

@heena , how do you know F= mv^2/r (Centripetal force) ?? @JerJason your question is value in nature as it does not explain the kind of force you are dealing with . There are many types of forces and for each force, there exists different relation with velocity,

OpenStudy (anonymous):

centrpetal force exerted by earth by that i think ...

OpenStudy (anonymous):

Well, the kind of force is not described in the question and hence we cannot comment on the relation between Force and Velocity. If the question pertains to Centripetal Force in specific, then what you have written is right :)

OpenStudy (anonymous):

The above comment was for @heena

OpenStudy (anonymous):

Sorry I did'nt have much time when I had originally made this post. But yes this is a question regarding centrifugal force.

OpenStudy (anonymous):

Yes, the above explanation by @heena is right :D Follow that @JerJason

OpenStudy (anonymous):

ok I will thanks.

OpenStudy (anonymous):

lol @JerJason atleat when u post qn plz write full qn :)

OpenStudy (anonymous):

Sure .